### v2

@@ 4.2.3 Formula Generation
To determine the best formulas for our data set, we divide the set of GRBs with $z_{obs}$ into a training and test set with an 80:20 split. A total of 16510 formulas were created from combinations of the seven variables chosen by LASSO (see Sec. 4.2.1). Each formula was tested by performing a 10-fold cross-validation (10fCV) with the training set 100 times. 10fCV is an ML method for assessing the performance of a model. It splits a given training set into 10 parts (folds). The model is trained on 9 folds and predicts on the 10th fold. This step is repeated until the model has predicted for all 10 folds, and then the results are assimilated, providing a measure of the model’s performance. We thus obtain cross-validated Pearson correlation coefficient ($r$) and root-mean-square (RMSE) values for each formula. Based on these scores, we select a subset of formulas that have achieved $r$ above 99.9% quantile and RMSE score below 2% quantile (see red points in the left panel of Fig. 6). The exact cutoff for $r$ is 0.553, and for RMSE is 1.145. This subset of formulas is used for obtaining predictions on the test set, which we had initially created (see right panel of Fig. 6). Finally, we select the best formulas based on this test set’s performance. Specifically, we pick three formulas that achieve the highest $r$, lowest median absolute deviation (MAD), and lowest RMSE.
Figure 6: Left panel: Plot showing the RMSE and Correlation scores for 16510 formulas. The red points show the formulas selected for analysis on the test set. These formulas achieve correlation and RMSE scores above the cut-offs mentioned at the top of the figure. Right panel: Performance of the selected subset of formulas on the test set. Each plus symbol here represents one formula selected from the left panel (red points). The red vertical and horizontal lines point to the formula that obtained the lowest RMSE and highest correlation on the test set.
@@
However, during our analysis, we discovered that the formulas determined via the above-mentioned methodology are sensitive to the particular training set selected. Thus, minimizing this strong dependence on the training set and selecting formulas that generalize to the full set of GRBs is important. To accomplish this, we follow the steps mentioned above; however, we randomize the GRBs used in the training and test sets while maintaining the 80:20 split ratio. We created 100 pairs of training and test sets and performed the subsequent steps for each pair. Furthermore, all the above-mentioned steps were repeated independently for GAM and GLM.
@@
In addition to testing the 16510 formulas with GAM and GLM, we also tested exclusively 148 formulas with GAM that have smoothing terms. We note here that even though these 148 formulas with smoothing were included in the 16510 formula set, they were not chosen during the analysis. However, it is essential to test these because in our previous work, (Dainotti et al., 2024c), smoothing formulas worked better, as they can capture the nonlinear relations between features. Thus, we decided to test them separately. These 148 formulas include smoothing terms applied to features individually and in pairs. The methodology described above was applied to these 148 formulas, and we obtained two formulas that work best with our data set. These were included in our final ensemble.
@@
Figure 7: The figure shows the GAM/GLM Formula Generator page of the web app. This module allows users to access the best formula for predicting redshift based on the training set provided and its predictors. The user can also manually set the cut-off for the correlation and RMSE for selecting the formula.
@@ 4.3 Outlier removal
Before performing any ML training, we need to remove outliers, as these can affect the model performance. We use the M-estimator technique to remove outliers (Huber, 1996, 1964). This is a robust regression method where the outliers are determined based on a robust loss function. Unlike ordinary least squares regression, which is sensitive to outliers, M-estimator assigns lower weights to extreme observations, making the results more robust. For the outlier removal, we selected the formula (Eq. 1), which obtained the highest counts for GLM. Using this formula, the M-estimator fits the 238 GRBs and assigns weights ranging from 0 to 1. GRBs that obtained weights less than 0.65 were eliminated. This specific weight of 0.65 was selected because it removes 5% of the data set as outliers. Thus, we are left with 226 GRBs, which are split again into training and test sets containing 180 and 46 GRBs, respectively. This training set of 180 GRBs is 20% larger than the training set used in Dainotti et al. (2024c).
@@ 4.4 The Superlearner
SuperLearner is an ensemble method that combines multiple ML models and leverages these models to obtain the best prediction. SuperLearner assigns to each model a weight based on its performance (larger weight for better performance), and the sum of all weights is 1, as indicated in the following formula:
$\sum{W_{i}}=1$
where $W_{i}$ is the weight for the $i^{th}$ model. The performance of the constituent models is determined via the 10fCV method, which is built into SuperLearner. The advantage of the SuperLearner is that its prediction is a weighted average of the predictions of its constituent models, with better-performing models obtaining larger weights. However, different ML models have their own individual implicit assumptions and priors, which can impact the redshift estimation. Here, we use the SuperLearner weights to select models whose implicit priors best work with our dataset. Models that obtain higher weights are included in the final nested 100 times 10fCV procedure, which provides the results presented in this paper. Furthermore, we are combining parametric (GLM), semi-parametric (GAM) and non-parametric (RF) into the final SuperLearner model, highlighting the benefits of an ensemble approach. A linear model like GLM enables higher performance for linear relations in the data, while GAM and RF enable better fit for non-linear relations. Thus, the models share features along with their individual features, enabling a performative ensemble ML model.
@@
Figure 8: The figure depicts the SuperLearner page of the web app, showing the data upload panel and options to choose model and data parameters.
@@ 4.5 Balance sampling
We aim to investigate how a more balanced sample of sources will affect the performance of our ML model. A balanced sample is a data set in which the response variable’s distribution is close to uniform. Such a distribution helps train the model over the full range of potential values for the response variable, thus helping it generalize to future samples. In this analysis, we deal with an unbalanced sample where the number of GRBs at $z>2.5$ (80) is fewer than the number of GRBs at $z<2.5$ (158). This phenomenon can be attributed to the Malmquist bias effect, which causes us to only detect the brightest astrophysical sources at high-$z$ due to detector thresholds. This effect prevents us from seeing a large population of dim sources at high-$z$. If we had a perfect satellite with extremely high sensitivity, we could populate the distribution at high-$z$. There are several methods to balance the sample: up-sampling, down-sampling, and balance sampling. For up-sampling, the original data remains unchanged, and duplicate samples of the minority class are added. In down-sampling, data is randomly selected to ensure that all classes are equally represented, matching the frequency of the minority class.
For up-sampling, multiple methods exist. We applied the Synthetic Minority Over-sampling Technique (SMOTE, Chawla et al. (2011); Dal Pozzolo et al. (2013)). SMOTE, as the name implies, creates synthetic copies of the minority class, thereby increasing its sample count. This is achieved by generating new data points within the feature space of the minority class dictated by the real nearest-neighbor data points. This prevents overfitting, which can happen when minority class samples are duplicated. It also eliminates the need to remove samples from the majority class, which may lead to information loss and reduced accuracy.
The redshift distribution of the upsampled data is shown in Fig. 9. This balanced sample also undergoes the same data treatment process as mentioned in Sec. 4.3.
@@
Figure 9: The redshift distribution obtained with the upsampling and the original redshift distribution are shown in solid and dashed lines, respectively.
@@ 4.6 Bias correction
Bias is defined as the difference between the observed and predicted response variables. ML models trained on an imbalanced training set are more biased towards the range containing more examples. In our sample of GRBs, the models tend to predict more toward lower redshift ranges as there are more GRBs in that range (see Sec. 4.5). Balanced sampling can mitigate this issue to an extent. However, we apply the Optimal Transport bias correction technique (Ferradans et al., 2013) to address this issue more directly. Some of us have already used it in Dainotti et al. (2024c, b); Narendra et al. (2022).
In this technique, first, the observed and predicted values of the response variable are sorted in ascending order. Next, a linear model is fitted between them and the predicted values are corrected by multiplying the slope ($A)$ and adding the intercept ($B$) of the linear fit. Furthermore, bias correction is performed separately in three redshift ranges: $20<z<$ 2, $3.52<z<$ 3.5, and $8.43.5<z<$ 8.4. These particular ranges were selected as they produced the best results for our particular ML model and dataset. The bias correction equation is as follows:
$Z_{C}=B+A*Z_{P}$
@@ 4.7 Monte Carlo analysis
With the MC simulations, we aim to explore the reason for the changes in our results compared to Dainotti et al. (2024c). The MC simulations also help us assess the degree of uncertainties in the prediction once the sample size is increased. Since in the procedure detailed in Dainotti et al. (2024c), we have used MICE and then run the Superlearner on the imputed data, we here perform MC simulations on the imputed data to keep the comparison more straightforward.
We perform simulations on the 7 predictors which are selected by LASSO, see Fig. 5. We simulate three sample sizes: 14, 28, and 42. This reflects, on average, how many GRBs with redshift and plateau will be added to the Swift catalog in one, two, and three years, respectively. This average was calculated based on the training set currently being used. Each sample size was generated 200 times with unique instances and combined with the training set of observed GRBs. For sampling, we used the SMOTE method (see Sec. 4.5), as this enables multivariate sampling of the data. Multivariate sampling is important as it maintains the correlations between the seven parameters involved, such as between $\log(T_{a})$ and $\log(F_{a})$ and $\log(Peakflux)$ (Dainotti et al., 2008). We simulated the samples so that the distributions of the variables resemble the current ones, thus we retain only the samples that pass the KS test. Each simulated sample was tested with 10fCV using the ensemble SuperLearner model described in Sec. 4.4. The results from this analysis are presented in Sec. 5.4.
4.8 Model Comparison with web-app
@@
Figure 10: This figure shows the various options and parameters available to a user of the web app when using the Model Comparision functionality.
@@ 4.9 Relative Importance of variables
The Relative Importance module of the web app evaluates the relative significance of the variables used for predicting the redshift. The process is initiated by importing data files the user provides. The ML model used for this module is the best model determined from our analysis in this publication (see Sec. 4.2.2). The relative importance of the predictors is decided based on this model’s performance. Users can also use the pre-trained models from the web app’s Model Comparison or Superlearner modules to obtain relative importance values for their data.
@@
Figure 11: Upper panel: The user interface of the Relative Influence module of the web app. Lower panel: Relative importance bar plot for the predictors used in this analysis.
@@ 4.10 Data Visualization with the web app
Figure 12: This figure shows the interface for the Data visualization module on the web app.
@@ 5 Results
5.1 Performance of the ML model
The results shown here are obtained after performing 10fCV 100 times on 180 GRBs using the SuperLearner model as described in Sec. 4.4. The correlation plot between $z_{obs}$ and $z_{pred}$ obtained from SuperLearner is shown in the upper left panel of Fig. 13. We achieve an $r$=0.646, RMSE=1.011, bias=0.14 and NMAD = 1.34. We apply the optimal transport bias correction (see Sec. 4.6) to these results to obtain the plot shown in the upper right panel of Fig. 13. Here, we obtain $r$=0.89, RMSE = 0.62, bias = 0.0047, and NMAD = 0.86. Sec. 6 discusses how these results compare with those presented in Dainotti et al. (2024c) and its implications.
@@ 5.1 Performance of the ML model
Figure 13: Results obtained from performing 100 times 10fCV using the SuperLearner model. The bottom and upper panels present results with and without balanced sampling, respectively. The upper left panel shows the result obtained from the SuperLearner model. The upper right panel shows the corresponding bias-corrected results. The bottom left panel shows the result obtained from the SuperLearner model when using a balanced sample. The bottom right panel shows the corresponding bias-corrected results.
@@ 5.2 Relative importance results
  • The third factor is the evolution of the metallicity of the progenitor’s ejecta. Heintz et al. (2023) found that the effective metallicity derived from the GRB spectra does evolve with redshift, and thus, it might have some impact on the $\log(NH)-z$ relation. We expect the progenitors of GRBs at high-$z$ to have lower metallicity than their low-$z$ counterparts. This can lead to the complex evolution of absorption of the GRB’s emission and, thus, significantly affect the $\log(NH)-z$ relation.
@@ 5.3 Estimating new redshifts in the generalization sample
As stated in Sec. 3, we have increased the generalization sample by 35%, from 221 GRBs (Dainotti et al., 2024c) to 276 GRBs. The histogram plots comparing the training and generalization sets’ features are shown in Fig. 14. We compared the distribution of each parameter between the generalization set and the training set via the KS test, and the p-values are stated at the top of each parameter distribution.
@@
Figure 15: This figure shows the scatter matrix plot of the generalization and training sets. The generalization set is presented with orange dots, and the training set is presented with cyan dots. The diagonal shows the histogram distribution for the generalization set, while the upper triangle shows the Pearson correlation values for respective pairs of parameters.
@@
We show a boxplot distribution of the predicted redshifts in the lower panel of Fig. 5.3. The red and grey boxplots show the bias-corrected and non-bias-corrected distributions. The x-axis shows the names of the generalization set GRBs, and the y-axis shows the redshift values. Note that because of the image’s scale, not all the GRB names are present along the axis. Our work here has enabled us to increase the Swift sample with pseudo-redshifts by 110%.
Figure 16: Upper panel: The redshift distribution of the generalization and training sets. The training set is shown with red histograms, while the generalization set with and without bias correction is shown in white and green histograms, respectively. Lower panel: This figure shows the box plot distribution of the generalization set redshift predictions, sorted from lowest to highest redshift. The bias-corrected and non-bias-corrected distributions are shown in red and grey boxplots, respectively. The dots show the outlier predictions obtained.

### v3

@@ 4.2.3 Formula generation
To determine the best formulas for our data set, we divided the set of GRBs with $z_{obs}$ into a training set and a test set with an 80:20 split. A total of 16510 formulas were created from combinations of the seven variables chosen by LASSO (see Sec. 4.2.1). Each formula was tested by performing a ten-fold cross-validation (10fCV) with the training set 100 times. The 10fCV is an ML method for assessing the performance of a model. It splits a given training set into ten parts (folds). The model is trained on nine folds and predicts on the tenth fold. This step is repeated until the model has predicted for all ten folds, and then the results are assimilated, providing a measure of the model’s performance. We thus obtained a cross-validated Pearson correlation coefficient ($r$) and root mean square error (RMSE) values for each formula. Based on these scores, we select a subset of formulas that have achieved $r$ above 99.9% quantile and RMSE score below 2% quantile (see red points in the left panel of Fig. 6). The exact cutoff for $r$ is 0.565, and for RMSE is 1.167. This subset of formulas is used for obtaining predictions on the test set, which we had initially created (see right panel of Fig. 6). Finally, we selected the best formulas based on this test set’s performance. Specifically, we picked three formulas that achieve the highest $r$, lowest median absolute deviation (MAD), and lowest RMSE:
Figure 6: Left panel: Plot showing the RMSE and correlation scores for 16510 formulas. The red points show the formulas selected for analysis on the test set. These formulas achieve correlation and RMSE scores above the cutoffs mentioned at the top of the figure. Right panel: Performance of the selected subset of formulas on the test set. Each plus symbol here represents one formula selected from the left panel (red points). The red vertical and horizontal lines point to the formula that obtained the lowest RMSE and highest correlation on the test set.
@@
However, during our analysis, we discovered that the formulas determined via the above-mentioned methodology are sensitive to the particular training set selected. Thus, minimizing this strong dependence on the training set and selecting formulas that generalize to the full set of GRBs is important. To accomplish this, we follow the steps mentioned above; however, we randomized the GRBs used in the training and test sets while maintaining the 80:20 split ratio. We created 100 pairs of training and test sets and performed the subsequent steps for each pair. Furthermore, all the above-mentioned steps were repeated independently for GAM and GLM.
@@
In addition to testing the 16510 formulas with GAM and GLM, we also tested exclusively 148 formulas with GAM that have smoothing terms. We note here that even though these 148 formulas with smoothing were included in the 16510 formula set, they were not chosen during the analysis. However, it is essential to test these because in our previous work, (Dainotti et al., 2024c), smoothing formulas worked better, as they can capture the non-linear relations between features. Thus, we decided to test them separately. These 148 formulas include smoothing terms applied to features individually and in pairs. The methodology described above was applied to these 148 formulas, and we obtained two formulas that work best with our data set. These were included in our final ensemble.
@@
Figure 7: Image of the GAM and GLM Formula Generator page of the web app. This module allows users to access the best formula for predicting redshift based on the training set provided and its predictors. The user can also manually set the cutoff for the correlation and RMSE in order to select the formula.
@@ 4.3 Outlier removal
Before performing any ML training, we need to remove outliers, as these can affect the model performance. We use the M-estimator technique to remove outliers (Huber, 1996, 1964). This is a robust regression method where the outliers are determined based on a robust loss function. Unlike ordinary least squares regression, which is sensitive to outliers, M-estimator assigns lower weights to extreme observations, making the results more robust. For the outlier removal, we selected the formula (Eq. 1), which obtained the highest counts for GLM. Using this formula, the M-estimator fits the 238 GRBs and assigns weights ranging from 0 to 1. GRBs that obtained weights less than 0.65 were eliminated. This specific weight of 0.65 was selected because it removes 5% of the data set as outliers. Thus, we were left with 226 GRBs, which were split again into training and test sets containing 180 and 46 GRBs, respectively. This training set of 180 GRBs is 20% larger than the training set used in Dainotti et al. (2024c).
@@ 4.4 The SuperLearner
SuperLearner is an ensemble method that combines multiple ML models and leverages these models to obtain the best prediction. SuperLearner assigns to each model a weight based on its performance (larger weight for better performance), and the sum of all weights is one, as indicated in the following formula:
$\sum{W_{i}}=1.$
Here, $W_{i}$ is the weight for the $i^{th}$ model. The performance of the constituent models is determined via the 10fCV method, which is built into SuperLearner. The advantage of the SuperLearner is that its prediction is a weighted average of the predictions of its constituent models, with better-performing models obtaining larger weights. However, different ML models have their own individual implicit assumptions and priors, which can impact the redshift estimation. Here, we use the SuperLearner weights to select models whose implicit priors best work with our data set. Models that obtain higher weights are included in the final nested 100 times 10fCV procedure, which provides the results presented in this paper. Furthermore, we are combining parametric (GLM), semi-parametric (GAM) and non-parametric (RF) into the final SuperLearner model, highlighting the benefits of an ensemble approach. A linear model such as GLM enables higher performance for linear relations in the data, while GAM and RF enable a better fit for non-linear relations. Thus, the models share features along with their individual features, enabling a performative ensemble ML model.
@@
Figure 8: Image of the SuperLearner page of the web app. Shown are the data upload panel and options to choose model and data parameters.
@@ 4.5 Balance sampling
We aim to investigate how a more balanced sample of sources will affect the performance of our ML model. A balanced sample is a data set in which the response variable’s distribution is close to uniform. Such a distribution helps train the model over the full range of potential values for the response variable, thus helping it generalize to future samples. In this analysis, we deal with an unbalanced sample where the number of GRBs at $z>2.5$ (80) is fewer than the number of GRBs at $z<2.5$ (158). This phenomenon can be attributed to the Malmquist bias effect, which causes us to only detect the brightest astrophysical sources at high-$z$ due to detector thresholds. This effect prevents us from seeing a large population of dim sources at high-$z$. If we had a perfect satellite with extremely high sensitivity, we could populate the distribution at high-$z$. There are several methods to balance the sample: up-sampling, down-sampling, and balance sampling. For up-sampling, the original data remains unchanged, and duplicate samples of the minority class are added. In down-sampling, data are randomly selected to ensure that all classes are equally represented, matching the frequency of the minority class.
For up-sampling, multiple methods exist. We applied the Synthetic Minority Over-sampling Technique (SMOTE, Chawla et al. (2011); Dal Pozzolo et al. (2013)). As the name implies, SMOTE creates synthetic copies of the minority class, thereby increasing its sample count. This is achieved by generating new data points within the feature space of the minority class dictated by the real nearest-neighbor data points. This prevents overfitting, which can happen when minority class samples are duplicated. It also eliminates the need to remove samples from the majority class, which may lead to information loss and reduced accuracy.
The redshift distribution of the upsampled data are shown in Fig. 9. This balanced sample also undergoes the same data treatment process as mentioned in Sec. 4.3.
@@
Figure 9: Redshift distribution obtained with the upsampling (solid lines) and the original redshift distribution (dashed lines).
@@ 4.6 Bias correction
Bias is defined as the difference between the observed and predicted response variables. ML models trained on an imbalanced training set are more biased toward the range containing more examples. In our sample of GRBs, the models tend to predict more toward lower redshift ranges as there are more GRBs in that range (see Sec. 4.5). Balanced sampling can mitigate this issue to an extent. However, we applied the optimal transport bias correction technique (Ferradans et al., 2013) to address this issue more directly. Some of us have already used it in Dainotti et al. (2024c, b); Narendra et al. (2022).
In this technique, first, the observed and predicted values of the response variable are sorted in ascending order. Next, a linear model is fit between them and the predicted values are corrected by multiplying the slope ($A)$ and adding the intercept ($B$) of the linear fit. Furthermore, bias correction is performed separately in three redshift ranges: $20<z<$ 2, $3.52<z<$ 3.5, and $8.43.5<z<$ 8.4. These particular ranges were selected as they produced the best results for our particular ML model and data set. The bias correction equation is as follows:
$Z_{C}=B+A*Z_{P},$
@@ 4.7 Monte Carlo analysis
With the MC simulations, we aim to explore the reason for the changes in our results compared to Dainotti et al. (2024c). The MC simulations also help us assess the degree of uncertainties in the prediction once the sample size is increased. In the procedure detailed in Dainotti et al. (2024c), we applied MICE to impute the missing data and then ran Superlearner on the imputed data. Therefore, in this work also we perform MC simulations on the imputed data to keep the comparison more straightforward.
We perform simulations on the 7 predictors which are selected by LASSO, see Fig. 5. We simulate three sample sizes: 14, 28, and 42. This reflects, on average, how many GRBs with redshift and plateau will be added to the Swift catalog in one, two, and three years, respectively. This average was calculated based on the training set currently being used. Each sample size was generated 200 times with unique instances and combined with the training set of observed GRBs. For sampling, we used the SMOTE method (see Sec. 4.5), as this enables multivariate sampling of the data. Multivariate sampling is important as it maintains the correlations between the seven parameters involved, such as between $\log(T_{a})$ and $\log(F_{a})$ and $\log(Peakflux)$ (Dainotti et al., 2008). We simulated the samples so that the distributions of the variables resemble the current ones, and thus we retained only the samples that pass the KS test. Each simulated sample was tested with 10fCV using the ensemble SuperLearner model. The results from this analysis are presented in Sec. 5.4.
4.8 Model comparison with web app
@@
Figure 10: Various options and parameters available to a user of the web app when using the Model Comparison functionality.
@@ 4.9 Relative importance of variables
The Relative Importance module of the web app evaluates the relative significance of the variables used for predicting the redshift. The process is initiated by importing data files the user provides. The ML model used for this module is the best model determined from our analysis in this publication (see Sec. 4.2.2). The relative importance of the predictors is decided based on this model’s performance. Users can also use the pretrained models from the web app’s Model Comparison or Superlearner modules to obtain relative importance values for their data.
@@
Figure 11: Upper panel: User interface of the Relative Influence module of the web app. Lower panel: Relative importance bar plot for the predictors used in this analysis.
@@ 4.10 Data visualization with the web app
Figure 12: Interface for the data visualization module of the web app.
@@ 5 Results
5.1 Performance of the machine learning model
The results shown here are obtained after performing 10fCV 100 times on 180 GRBs using the SuperLearner model. The correlation plot between $z_{obs}$ and $z_{pred}$ obtained from SuperLearner is shown in the upper left panel of Fig. 13. We achieve an $r$=0.646, RMSE=1.011, bias=0.14 and NMAD = 1.34. We apply the optimal transport bias correction (see Sec. 4.6) to these results to obtain the plot shown in the upper right panel of Fig. 13. Here, we obtain $r$=0.89, RMSE = 0.62, bias = 0.0047, and NMAD = 0.86. Sec. 6 discusses how these results compare with those presented in Dainotti et al. (2024c) and its implications.
@@ 5.1 Performance of the machine learning model
Figure 13: Results obtained from performing 10fCV 100 times using the SuperLearner model. The bottom and upper panels present results with and without balanced sampling, respectively. The upper-left panel shows the results obtained from the SuperLearner model. The upper-right panel shows the corresponding bias-corrected results. The bottom-left panel shows the result obtained from the SuperLearner model when using a balanced sample. The bottom-right panel shows the corresponding bias-corrected results.
@@ 5.2 Relative importance results
  • The third factor is the evolution of the metallicity of the progenitor’s ejecta. Heintz et al. (2023) found that the effective metallicity derived from the GRB spectra does evolve with redshift, and thus, it might have some impact on the $\log(NH)-z$ relation. We expect the progenitors of GRBs at high-$z$ to have lower metallicity than their low-$z$ counterparts. This can lead to the complex evolution of absorption of the GRB’s emission and thus significantly affect the $\log(NH)-z$ relation.
@@ 5.3 Estimating new redshifts in the generalization sample
As stated in Sec. 3, we have increased the generalization sample by 35%, from 221 GRBs (Dainotti et al., 2024c) to 276 GRBs. The histogram and scatter matrix plots comparing the training and generalization sets’ features are shown in Fig. 14 and Fig. 15, respectively. We compared the distribution of each parameter between the generalization set and the training set via the KS test, and the p-values are stated at the top of each parameter distribution.
@@
Figure 15: Scatter matrix plot of the generalization and training sets. The generalization set is presented with orange dots, and the training set is presented with cyan dots. The diagonal shows the histogram distribution for the generalization set, while the upper triangle shows the Pearson correlation values for respective pairs of parameters.
@@
We show a box plot distribution of the predicted redshifts in the lower panel of Fig. 5.3. The red and gray box plots show the bias-corrected and non-bias-corrected distributions. The x-axis shows the names of the generalization set GRBs, and the y-axis shows the redshift values. We note that because of the image’s scale, not all the GRB names are present along the axis. Our work here has enabled us to increase the Swift sample with pseudo-redshifts by 110%.
Figure 16: Upper panel: Redshift distribution of the generalization and training sets. The training set is shown with red histograms, while the generalization set with and without bias correction is shown in white and green histograms, respectively. Lower panel: Box plot distribution of the generalization set redshift predictions, sorted from lowest to highest redshift. The bias-corrected and non-bias-corrected distributions are shown in red and gray box plots, respectively. The dots show the outlier predictions obtained.